# TRABSA: Interpretable Sentiment Analysis of Tweets using Attention-based BiLSTM and Twitter-RoBERTa

### v2

@@ 3 Hidden Layers NN 94% 93% 93% 94%
The LIME method provides a methodical technique for evaluating individual predictions given by complicated ML models, which stands for Locally Interpretable Model-Agnostic Explanations [43]. This approach works by estimating the model’s behavior around a given forecast. Fundamentally, LIME utilizes a local linear explanation model that follows Equation 9 to the letter, making it an additive feature attribution method. LIME presents the idea of “interpretable inputs," which are condensed representations of the original inputs and are represented as $x_{0}$. A binary vector of interpretable inputs is mapped to the original input space via the transformation $x=h_{x}(x_{0})$. Different kinds of mappings $h_{x}$ are used for different input spaces. For example, when applied to bag-of-words text characteristics, $h_{x}$ translates a binary vector (signaling the presence or absence of words) to the appropriate word count in the source text. LIME aims to minimize the objective function to determine the coefficients $\phi$:
@@
Our methodology and results demonstrate a significant advancement in sentiment analysis compared to existing literature. While prior studies have explored diverse models and techniques, our TRABSA model introduces a unique hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. This innovative combination enables our model to effectively capture nuanced sentiment patterns, resulting in notably higher accuracy and performance across multiple evaluation metrics than traditional and state-of-the-art transformer models such as BERT and RoBERTa. We conducted a thorough analysis to assess the robustness of our TRABSA model across various datasets and scenarios, consistently observing superior performance across multiple datasets, including extended and external ones. Additionally, our model demonstrates resilience to variations in sentiment expression and context, reaffirming its reliability in diverse real-world scenarios.
This novel hybrid approach offers several benefits to the field, including unparalleled accuracy, robustness, and generalizability across diverse datasets and scenarios. By leveraging the strengths of each component, the TRABSA model can revolutionize sentiment analysis applications, providing researchers, businesses, and policymakers with deeper insights into public opinion, consumer sentiment, and social trends. Its innovative architecture and superior performance represent a significant advancement in the quest for more accurate and reliable sentiment analysis tools, with implications extending beyond academic research.
@@
Our research has yielded significant findings and contributions to sentiment analysis. We have achieved remarkable results by developing and evaluating the TRABSA model, a novel hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. Leveraging the latest RoBERTa-based transformer model and expanding the datasets, we have demonstrated the TRABSA model’s exceptional accuracy and relevance, bridging existing gaps in sentiment analysis benchmarks. Thorough comparisons of word embedding techniques and methodical labeling of tweets using lexicon-based approaches have further enhanced the effectiveness of sentiment analysis methodologies. Our experiments and benchmarking efforts have highlighted the superiority of the TRABSA model over traditional and state-of-the-art models, showcasing its versatility and robustness across diverse datasets and scenarios. With macro-average precision of 94%, macro-average recall of 93%, macro-average F1-score of 94%, and accuracy of 94%, our model has proven its efficacy in capturing nuanced sentiment patterns. Additionally, exploring model interpretability techniques using SHAP and LIME has enhanced our understanding and trust in the TRABSA model’s predictions, reinforcing its practical applicability.
Despite the significant advancements achieved in our research, several avenues remain for future exploration and improvement in interpretable sentiment analysis. Firstly, there is scope for refining and expanding model interpretability techniques to provide deeper insights into the factors influencing sentiment predictions. Additionally, integrating multimodal data sources such as text, images, and audio could enhance the richness and accuracy of sentiment analysis. Addressing ethical considerations regarding bias, fairness, and privacy in sentiment analysis models is paramount for responsible deployment and usage. Furthermore, exploring the application of sentiment analysis in emerging domains such as healthcare, finance, and politics could uncover new challenges and opportunities for research and innovation. Overall, continued research in interpretable sentiment analysis holds the potential to drive meaningful advancements in AI technologies and contribute to more informed decision-making in various fields.
@@
(https://www.kaggle.com/datasets/crowdflower/twitter-airline-sentiment).

### v3

@@ 3 Hidden Layers NN 94%$\pm$293%$\pm$393%$\pm$394%$\pm$21166 53
The LIME method provides a methodical technique for evaluating individual predictions given by complicated ML models, which stands for Locally Interpretable Model-Agnostic Explanations [60]. This approach works by estimating the model’s behavior around a given forecast. Fundamentally, LIME utilizes a local linear explanation model that follows Equation 9 to the letter, making it an additive feature attribution method. LIME presents the idea of "interpretable inputs," which are condensed representations of the original inputs and are represented as $x_{0}$. A binary vector of interpretable inputs is mapped to the original input space via the transformation $x=h_{x}(x_{0})$. Different kinds of mappings $h_{x}$ are used for different input spaces. For example, when applied to bag-of-words text characteristics, $h_{x}$ translates a binary vector (signaling the presence or absence of words) to the appropriate word count in the source text. LIME aims to minimize the objective function to determine the coefficients $\phi$:
@@
Our methodology and results demonstrate a significant advancement in SA compared to existing literature. While prior studies have explored diverse models and techniques, our TRABSA model introduces a unique hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. This innovative combination enables our model to effectively capture nuanced sentiment patterns, resulting in notably higher accuracy and performance across multiple evaluation metrics than traditional and state-of-the-art transformer models such as BERT and RoBERTa. We conducted a thorough analysis to assess the robustness of our TRABSA model across various datasets and scenarios, consistently observing superior performance across multiple datasets, including extended and external ones. Additionally, our model demonstrates resilience to variations in sentiment expression and context, reaffirming its reliability in diverse real-world scenarios.
This novel hybrid approach offers several benefits to the field, including unparalleled accuracy, robustness, and generalizability across diverse datasets and scenarios. By leveraging the strengths of each component, the TRABSA model can revolutionize SA applications, providing researchers, businesses, and policymakers with deeper insights into public opinion, consumer sentiment, and social trends. Its innovative architecture and superior performance represent a significant advancement in the quest for more accurate and reliable SA tools, with implications extending beyond academic research.
@@
Our research has yielded significant findings and contributions to SA. We have achieved remarkable results by developing and evaluating the TRABSA model, a novel hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. Leveraging the latest RoBERTa-based transformer model and expanding the datasets, we have demonstrated the TRABSA model’s exceptional accuracy and relevance, bridging existing gaps in SA benchmarks. Thorough comparisons of word embedding techniques and methodical labeling of tweets using lexicon-based approaches have further enhanced the effectiveness of SA methodologies. Our experiments and benchmarking efforts have highlighted the superiority of the TRABSA model over traditional and state-of-the-art models, showcasing its versatility and robustness across diverse datasets and scenarios. With macro-average precision of 94%, macro-average recall of 93%, macro-average F1-score of 94%, and accuracy of 94%, our model has proven its efficacy in capturing nuanced sentiment patterns. Additionally, exploring model interpretability techniques using SHAP and LIME has enhanced our understanding and trust in the TRABSA model’s predictions, reinforcing its practical applicability.
Despite the significant advancements achieved in our research, several avenues remain for future exploration and improvement in interpretable SA. Firstly, there is scope for refining and expanding model interpretability techniques to provide deeper insights into the factors influencing sentiment predictions. Additionally, integrating multimodal data sources such as text, images, and audio could enhance the richness and accuracy of SA. Addressing ethical considerations regarding bias, fairness, and privacy in SA models is paramount for responsible deployment and usage. Furthermore, exploring the application of SA in emerging domains such as healthcare, finance, and politics could uncover new challenges and opportunities for research and innovation. Overall, continued research in interpretable SA holds the potential to drive meaningful advancements in AI technologies and contribute to more informed decision-making in various fields.
@@
(https://www.kaggle.com/datasets/crowdflower/twitter-airline-sentiment). The code for reproducibility is available in
https://github.com/Abrar2652/nlp-roBERTa-biLSTM-attention.